OpenStudy (anonymous):

solve the inequality (1/x)-x>0 I know the anwer is 0

OpenStudy (anonymous):

1 - x^2 > 0 x^2 < 1 (x+1)(x-1) < 0 so 0 < x < 1

OpenStudy (anonymous):

did you get this?

OpenStudy (anonymous):

but wouldn't you also get x<-1?

OpenStudy (anonymous):

no I didn't

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

first of all you have to put the left hand side of the inequality as one fraction

OpenStudy (anonymous):

oh yeah i was guessing im not sure of the answer

OpenStudy (anonymous):

okay so we have a common denominator....

OpenStudy (anonymous):

i screwed that up royally

OpenStudy (anonymous):

\[\frac{1}{x}-x>0\] \[\frac{1-x^2}{x}>0\]

OpenStudy (anonymous):

that is the right way to add sorry

OpenStudy (anonymous):

now factor the numerator, and you get \[\frac{(1+x)(1-x)}{x}>0\]

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so we have 3 factors to consider \[1+x\] \[x\] \[1-x\] and they are 0 at the numbers \[-1,0,1\] respectively

OpenStudy (anonymous):

alright with you soo far

OpenStudy (anonymous):

so we have to break up the real line into 4 different parts like this ___________-1_____________0______________1____________ and consider what happens on each part

OpenStudy (anonymous):

perhaps the easiest thing to do in this case is make a sign chart for each of the factors. \[1+x\] is zero at -1, positive for bigger numbers, negative for smaller. so i will write \[1+x\] ------------------- -1 ++++++++++++++++++++++++++ so indicate is is negative to the left and positive to the right

OpenStudy (anonymous):

\[x\] is evidently positive if x is positive (who is buried in grant's tomb) and negative if x is negative so the sign chart will look like \[x\] -------------------------------------- 0 ++++++++++++++++++++++ indicating it is positive to the left and negative to the right

OpenStudy (anonymous):

\[1-x\] is the tricky one. it is 0 if x is 1, but it is positive to the RIGHT AND NEGATIVE TO THE LEFT of 1 so its sign chart will look like \[1-x\] ++++++++++++++++++++++++++++++++ 1 ------------------------

OpenStudy (anonymous):

now we pile them one on top of the other and get ------------------- -1 ++++++++++++++++++++++++++++++++++ --------------------------------------- 0 +++++++++++++++++++++++ +++++++++++++++++++++++++++++++++++++ 1 -------------------

OpenStudy (anonymous):

if i have lost you i understand but i cannot think of a snap way to do this now we just look. if \[x<-1\] two factors are negative, one is positive so answer is positive

OpenStudy (anonymous):

yeah it would be positive between 0 and 1 i got you!

OpenStudy (anonymous):

so solution is... \[x<-1\] or \[0<x<1\]

OpenStudy (anonymous):

tada!

OpenStudy (anonymous):

one quick Q

OpenStudy (anonymous):

whew ok hope that one was clear

OpenStudy (anonymous):

if you tried to make it a common denomintor wouldn't you have to multiply everything with x making it 1-x^/x not the way u wrote it?

OpenStudy (anonymous):

i hope that is what i wrote \[\frac{1}{x}-x=\frac{1}{x}-\frac{x^2}{x}=\frac{1-x^2}{x}\]

OpenStudy (anonymous):

let me scroll up and look

OpenStudy (anonymous):

yeah it is there. hope it is clear the problem with these is that you are not free to multiply one side or the other by x because you don't know if x is positive or negative. so they require more work

OpenStudy (anonymous):

no Im saying that when you multiplied the -x by 1/x dont you have to multiply the other parts by x too?

OpenStudy (anonymous):

no \[\frac{1}{x}\] is already over x

OpenStudy (anonymous):

forget variables, let's do it with numbers \[\frac{1}{3}-3\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you have to take \[\frac{1-9}{3}\] right?

OpenStudy (anonymous):

that is you need \[\frac{1}{3}-\frac{9}{3}\] in order to subtract

OpenStudy (anonymous):

oh okay that makes sense thanks :)

OpenStudy (anonymous):

yw